Symmetry: Culture and Science Vol. 21, Nos. 1-4, 333-336, 2010

SECTION TITLE

TITLE OF THE ARTICLE: WITH CAPITAL LETTERS

Firstname Authorlastnam1 ^{1,∗}, Firstname Authorlastnam2 ^{2}

1 Optional data: Mathematician, (b. Boston, Mass., USA, 1938). Address: Professional or personal address of author 1 (required). E-mail: email.address@author1. Optional URL. Next biographic entries are optional. Fields of interest: Geometry, mathematical crystallography (also ornamental arts, anthropology - non-professional interests in parentheses). Awards: Symmetry Award, 1987; Dissymmetry Medal, 1989. Publications and/or Exhibitions: List your main symmetry-related publications/exhibitions in chronological order, following the conventions of the references. Please, give no more than the five most characteristic items here. 2 Address: Professional or personal address of author 2 (required). E-mail: email.address@author2. Optional URL. Optional biographic entries, as for author 1.

∗ Corresponding author

Abstract: The abstract of your paper in italic. Please summarize here what is original in the paper: methods, results, etc. The abstract should summarize in a few sentences what the readers and the reviewers can find in your paper. Please think that several databases may incorporate your abstract without change, and that your paper can be retrieved from it.

Keywords: keyword1, keyword2, keyword3, other keywords. PACS: 11.30.Qc, 24.80.+y, 33.15.Bh, 61.50.Ah, 87.15.B-, 78.20.Ek MSC: 11G42, 14J33, 52B15, 53D37, 58D19, 58J70, 76M60, 81R40

1. INTRODUCTION

Introduction, state of the art, etc. Please realize that the readers of the Symmetry Journal are not necessarily experts in your field. Please cite exact references.

334  A.B. Authorlastnam1, C.D. Authorlastnam2

2. THEORY

Methods. General paragraph. General paragraph. Noindentation. 2.1 Subsection Optional subsection here. 2.1.1 Sub-subsection Optional sub-subsection here. 2.2 Subsection An other optional subsection here.

3. RESULTS

Results. 3.1 Subsection Optional subsection here.

TABLE HERE (CENTERED OR NOT)

Table 1: legend of table 1

Main text. Comments.

HIGH DEFINITION FIGURE HERE (CENTERED OR NOT)

Figure 1: legend of figure 1

Main text. Comments.

Running title of the Article  335

3.2 Subsection Optional subsection here.

ANOTHER HIGH DEFINITION FIGURE HERE (CENTERED OR NOT)

Figure 2: legend of figure 2

Main text. Comments.

4. DISCUSSION

Discussion. 4.1 Subsection Optional subsection here.

5. CONCLUSION

Conclusion.

Acknowledgements

If you wish, you may express thanks to any colleague, peer, etc., who contributed to your paper with advices, ideas, comments, corrections or any concrete professional or technical assistance. Acknowledgement to the reviewers should be avoided, (unless they contributed essentially to change the scholarly content of the paper). Please avoid titles like Prof., Dr., etc.

REFERENCES

Asymmetrist, A.Z. (or corporate author) (1981) Book Title: Subtitle. Series Title no. 27, 2nd ed., City (only the first one): Publisher, vii + 619 pp. (further data can be added, e.g.: 3rd ed., 2 vols., ibid., 1985, viii + 444 + 484 pp. with 2 computer diskettes; Reprint, ibid., 1988; German trans., German Title. 2 vols., City: Publisher, 1990, 986 pp.; Hungarian trans.).

Asymmetrist, A.Z., Dissymmetrist, S.Z., and Symmetrist, W.M. (1980) Article or e-mail article title: Subtitle. Parts 1-2, Journal Name Without Abbreviation, [E-Journal or Discussion Group address: http://domain1.domain2.dom (if applicable)], 22 (volume number), no. 6 (issue number if each one restarts pagination), 110-119 (page numbers); 23, no. 1, 117-132 and 133 (for e-journals any appropriate data).

Dissymmetrist, S.Z. (1989a) Chapter, article, symposium paper, or abstract title. [Abstract (if applicable)], In: Editorologist, A.B. and Editorologist, C.D., eds. Book, Special Issue, Proceedings, or Abstract Volume Title, [Special Issue (or) Congress organized by the Dissymmetry Society, University of Symmetry, San Symmetrino, Calif., December 11-22, 1971 (those data which are not available from the title, if applicable)], Vol. 2, City: Publisher, 19-20 (for special issues the data of the journal).

Dissymmetrist, S.Z. (1989b) Dissertation Title. [Ph.D. Dissertation], City: Institution, 248 pp. (Exhibition Catalogs, Manuscripts, Master’s Theses, Mimeographs, Patents, Preprints, Working Papers, etc. in a similar way; Audiocassettes, Audiotapes, Compact Disks, Computer Diskettes, Computer Software, Films, Microfiches, Microfilms, Slides, Sound Disks, Videocasettes, etc. with necessary modifications, adding the appropriate technical data).

Dissymmetrist, S.Z., ed. (1990) Dissimmetriya v nauke (title in original, or transliterated, form). [Dissymmetry in science, in Russian with German summary], Trans. from English by B.W. Antisymmetrist, etc.

Phyllotaxist, F.B. (1899/1972) Title of the 1972 Edition. [Reprint, or Translation, of the 1899

ed.], etc. [Symmetrist, W.M.] (1989) Review of Title of the Reviewed Work, by S.Z. Dissymmetrist. etc.

(if the review has an additional title, then it should appear first; if the authorship of a work is not revealed in the publication, but known from other sources, the name should be enclosed in brackets).

Asymmetrist, A.Z. and Dissymmetrist, S.Z., eds. (1981) Book Title: Subtitle. Series Title no. 27, 2nd ed., City (only one): Publisher, vii + 619 pp. (further data can be added, e.g.: 3rd ed., 2 Vols., ibid., 1985, viii + 444 + 484 pp. and two computer diskettes; Reprint, ibid., German trans., German title. 2 Vols., City: Publisher, 1990, 986 pp.; Hungarian trans.)

Asymmetrist, A.Z., Dissymmetrist, S.Z., and Symmetrist, W.M. (1980-81) Article title: Subtitle. Parts 1-3, Journal Name Without Abbreviation, 22, (volume number) no. 5 (issue number if each one restarts pagination), 110-119 (page numbers); no. 6, 17-32; 23, no. 1, 47-74 and 77. (In case of E-mail article: Electronic Journal or Discussion Group address, date)

Dissymmetrist, S.Z. (1988a) Chapter or article title. In: Harmonist, B.A.C., ed., Book Title, City: Publisher, 112-123 (or) Special Issue Title, [Special Issue], (or) Journal Title Without Abbreviation, 18, nos. 1-2, 112-123.

Dissymmetrist, S.Z. (1988b) Dissimmetriya v nauke (title in original, or transliterated, form).[Dissymmetry in Science, in Russian with German summary], Trans. from English by B.W. Antisymmetrist, etc.

Symmetrist, W.M. (1972) Symposium paper or abstract title. [Abstract] (if applicable), In: Symposionist, A.B., ed. Proceedings or Abstract Volume Title, [Symposium organized by the Dissymmetry Society, University of Symmetry, San Symmetrino, Calif., December 11-22, 1971 (those data which are not available from the title)], Vol. 2, City: Publisher, 19 (page number).